OpenStudy (mortonsalt):

Applying Vector Differential Equations to Electrical Networks question! Will posted below.

OpenStudy (mortonsalt):

A 5V source has been connected in parallel with the capacitor in a series RLC circuit (R= 10 Ohms, C=2F, L=3H) for a long time. At t=0, the source is disconnected. Find a vector differential equation in matrix form which characterises the circuit (without the DC source) and determine the initial conditions.

OpenStudy (mortonsalt):

So I know that: \[i_c = C \frac{ dv_C }{ dt } + i_c(0^-)\] \[v_L = L\frac{di_l}{dt} + v_L(0^-)\]

OpenStudy (mortonsalt):

Through Kirchoff's Voltage Law, I know that \[v_c +iR + L\frac{di}{dt}=0\]

OpenStudy (mortonsalt):

And the relationship in matrix form should look something like \[\vec{x} = A\vec{x}\]

OpenStudy (mortonsalt):

sorry meant to put: \[\vec{x}'\]

OpenStudy (mortonsalt):

And from that point on, I'm not entirely sure what to do. :( Any help would be great, thanks!

ganeshie8 (ganeshie8):

\(v_c +iR + L\frac{di}{dt}=0\) replace \(i\) by \(dq/dt\) and \(v_c\) by \(q/C\) and you will see a DE in one variable \(q\) : \[L\dfrac{d^2q}{dt^2} + R\dfrac{dq}{dt}+\dfrac{1}{C}q=0\]

ganeshie8 (ganeshie8):

you can split that as a system of two differential equations : \(\dfrac{dq}{dt} = w\) \(\dfrac{dw}{dt} = -\dfrac{1}{LC}q-\dfrac{R}{L}w\)

ganeshie8 (ganeshie8):

see if you can write that in matrix form

ILovePuppiesLol (ilovepuppieslol):

i don't think she is here @ganeshie8

OpenStudy (mortonsalt):

@ILovePuppiesLol Still here! @ganeshie8 so what I did was I let \[\vec{x} = \left(\begin{matrix}v_c \\ i_L\end{matrix}\right)\] so \[i_c = x_2 = C\frac{dv_c}{dt}\] \[x_1 + x_2R + L\frac{dx_2}{dt}=0\] \[\frac{-x_1}L - \frac{x_2R}{L} = \frac{dx_2}{dt}\] \[\frac{dx_1}{dt} = \frac{x_2}{C}\] \[\frac{dx_2}{dt} = \frac{-x_1}{L}-\frac{x_2R}{L}\]

OpenStudy (mortonsalt):

So then the matrix is: \[\left[\begin{matrix}0 & 1/C \\ \frac{-1}{L} & \frac{-v(0^-)}{L}\end{matrix}\right]\left(\begin{matrix}v_c \\ i_L\end{matrix}\right)=\left(\begin{matrix}\frac{dv_c}{dt} \\ \frac{di_L}{dt}\end{matrix}\right)\]

OpenStudy (mortonsalt):

Then substitute the values in.

OpenStudy (mortonsalt):

Does this make sense? @ganeshie8

ganeshie8 (ganeshie8):

Looks good to me !